A company purchased new furniture at a cost of $16,000 on January 1. The furniture is estimated to have a useful life of 6 years and a $1,000 salvage value. The company uses the straight-line method of depreciation. What is the book value of the furniture on December 31 of the first year? Multiple Choice $13,500 $2,500 $15,000 $13,333 $16,000
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(16,000-1000)/6=$2500/year

Hence book value on December 31 of the first year=Cost-Accumulated Depreciation

=16,000-2500

=$13500.